I need help with two of these problems

Find a matrix P that orthogonally diagonalizes A = -1 201 2 02, then find PT AP. 0 2 1]

3 a alw /10 Given that A is orthogonal, what is the value of a? 1 3 V10 V10
Answer #1

2nd Giren that a mys A= 3 det (A) it A is orthogonal -0. Then , CO. NTO det(A) = a. the con la ate 9 wro Ando aa no 1st ! dodrectors, Now gen we have to find the corresponding ei for a=0 (A-T)* -0. 17 do a 2 -2 Ng 0 x + ad -zy 2X + 22 EO. + 2 = 0 zyfor 93-3. A-AI) x = 0. 2 2 76-6 2 2 4 3 o 2 22 t try By + 27 24 t 2y + 4z=e. z = old y=-e ket x=C 1 X= C WS an 2 Hence, p= [